OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a14 = -33, a15 = 9

OpenStudy (lasttccasey):

Assuming the difference is the pattern, 42 is added to -33 to get 9 so 9 + 42 gives...\[a_{16} = 51\]

OpenStudy (anonymous):

an = 4 • -3^n + 1 an = -3 • 4^n - 1 an = -3 • 4^n an = 4 • -3^n which of these will be the equation needed?

OpenStudy (lasttccasey):

None of those make sense.. any number to the 15th or 16th power will be massive.. Are you missing parenthesis or something?

OpenStudy (anonymous):

i am so sorry i mixed the answers up here are the correct options an = -579 + 42(n + 1) an = -579 + 42(n - 1) an = -579 - 42(n + 1) an = -579 - 42(n - 1)

OpenStudy (mathstudent55):

Plug in 14 and 15 into the choices, and see which one gives you a14 = -33 and a15 = 9.